# A solution of the problem of standard compact Clifford-Klein forms

### v1

@@ 4 Proof of Theorem 2: eliminating the regular case
We have $\tilde{Y}\notin\mathfrak{h}.$
@@
Since $x_{\alpha}+\sigma(x_{\alpha})\in\mathfrak{n}$ therefore there exist $H\in\mathfrak{n}_{h}$ and $L\in\mathfrak{n}_{l}$ s.t. $x_{\alpha}+\sigma(x_{\alpha})=H+L$ and so $\tilde{Y}=H+\theta(H)+L+\theta(L).$ If $\tilde{Y}\in\mathfrak{h}$ then $L+\theta(L)\in\mathfrak{h}.$ But
@@
and thus $L+\theta(L)=H_{1}+\theta(H_{1})$ for some $H_{1}\in\mathfrak{n}_{h}.$ As $\mathfrak{n}\cap\theta(\mathfrak{n})=\{0\}$ therefore $L\in\mathfrak{n}_{h}$ and so $L=0.$ Thus $x_{\alpha}+\sigma(x_{\alpha})\in\mathfrak{n}_{h}$ and $\alpha|_{\mathfrak{a}_{h}}=0,$ a contradiction. ∎
It follows that there exists a non-zero $Y\in\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma}$ which is orthogonal (with respect to $B$) to $\mathfrak{h}.$ Since $\mathfrak{h}^{c}$ is regular we obtain the following
Corollary 4.
There exists a non-zero $Y\in\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma}$ which centralizes $\mathfrak{h}.$
Theorem 8.
Assume that neither $G/H$ nor $H$ is compact. In the setting of this section (there exists $iX\neq 0$ and $\mathfrak{h}^{c}$ is regular in $\mathfrak{g}^{c}$) the pair $(\mathfrak{g},\mathfrak{h})$ cannot induce standard Clifford-Klein form.
@@
Put
$Z_{p}:=\sum_{\alpha\in\Delta_{p}^{\pm}}\mathfrak{g}_{\alpha},\ \ Z_{n}:=\sum_{\alpha\in\Delta_{n}^{\pm}}\mathfrak{g}_{\alpha},\ \ O:=\sum_{\alpha\in\Delta_{0}^{\pm}}\mathfrak{g}_{\alpha}.$
As $[Y,iX]=[Y,\mathfrak{a}_{h}]=0$ therefore
$[Y,Z_{n}]\subset Z_{n},\ \ [Y,Z_{p}]\subset Z_{p},\ \ [Y,O]\subset O.$
But $Y$ centralizes $\mathfrak{h}^{c}$ and so for $1\leq i\leq k$ we have $[Y,S_{h}^{i}+O_{i}]=0.$ By Theorem 4 (1.) we have
$[Y,x_{\alpha_{i}}]=0,\ \ \alpha_{i}\in\Delta_{p}^{\pm}$
and thus (as $[Y,\sigma(x_{\alpha_{i}})]=0$) $[Y,x_{\alpha_{j}}]=0$ for any $\alpha_{j}\in\Delta_{n}^{\pm}.$ By Lemma 7 this implies $[Y,x_{\beta_{i}}]=0$ for every $\beta_{i}\in\Delta_{0}^{\pm}.$ Thus
$[Y,\mathfrak{n}]=[Y,\theta(\mathfrak{n})]=0$
and so $[Y,\mathfrak{a}]=0$ as $\mathfrak{a}\subset[\mathfrak{n},\theta(\mathfrak{n})].$ But now $[Y,\mathfrak{p}]=0$ and so $[Y,\mathfrak{k}]=0$ since $\mathfrak{k}=[\mathfrak{p},\mathfrak{p}].$ Therefore $[Y,\mathfrak{g}]=0$ and $Y\neq 0.$ But now $\mathfrak{g}$ cannot be simple, a contradiction. ∎
@@ 5 Completion of proof of Theorem 2: non-regular case
In view of Theorem 8 we assume that the embeddings $\mathfrak{h}^{c},\mathfrak{l}^{c}\hookrightarrow\mathfrak{g}^{c}$ are not regular. The proof is divided into two parts:
@@
    $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ are simple,
@@
    at least one of $\mathfrak{h}^{c}$ or $\mathfrak{l}^{c}$ is semisimple but not simple.
The proof of the main theorem will follow from Propositions 6, 8, 9 proved in this section. We begin with case 1) and consider embeddings of classical Lie algebras $\mathfrak{h}^{c}$ as linear representations, and, following Dynkin [14], divide these into two classes: irreducible and reducible. In what follows we use the notation from the paper [14]. We use tables of inclusions of simple classical Lie algebras from [14], however, we also reproduce parts of them as Tables 3,4,5 in Section 6.
5.1 Proof in the case when $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ are simple
5.1.1 Irreducible representations and simple classical subalgebras
We begin with the classification of maximal simple classical subalgebras in simple Lie algebras (see, [30], Chapter 6, Theorem 3.1). Following Dynkin [14] we introduce the following terminology. Let $\mathfrak{g}^{c}$ be a classical Lie algebra, that is $\mathfrak{g}^{c}=\mathfrak{sl}(V),\mathfrak{so}(V),\mathfrak{sp}(V)$ where $V$ is a finite-dimensional complex linear space. Note that by assumption, $\mathfrak{h}^{c}$ is also classical, so, $\mathfrak{h}^{c}=\mathfrak{sl}(U),\mathfrak{so}(U),\mathfrak{sp}(U)$ for some finite-dimensional complex linear space.
@@ 5.1.1 Irreducible representations and simple classical subalgebras
Theorem 9 (Dynkin, [30]).
The classification of maximal semisimple Lie subalgebras $\mathfrak{h}^{c}$ in simple complex Lie algebras $\mathfrak{g}^{c}$ of classical type falls into the classes:
@@
    $\mathfrak{h}^{c}$ is reducible, and then a) $\mathfrak{h}^{c}$ is maximal parabolic, b) if $\mathfrak{g}^{c}=\mathfrak{so}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate to $\mathfrak{so}(k,\mathbb{C})\oplus\mathfrak{so}(n-k,\mathbb{C})$; if $\mathfrak{g}^{c}=\mathfrak{sp}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate to $\mathfrak{sp}(k,\mathbb{C})\oplus\mathfrak{sp}(n-k,\mathbb{C})$, $0<k<n$;
@@
    $\mathfrak{h}$ is irreducible and non-simple, and then the following possibilities may occur: $\mathfrak{g}^{c}=\mathfrak{sl}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate to $\mathfrak{so}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C}),$ $n=st,s,t\geq 2$, if $\mathfrak{g}^{c}=\mathfrak{so}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate either to $\mathfrak{so}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})$, where $n=st,s,t\geq 3$, $s,t\not=4$, or to $\mathfrak{sp}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})$, where $n=st,s,t\geq 2$, if $\mathfrak{g}^{c}=\mathfrak{sp}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate to $\mathfrak{sp}(s,\mathbb{C})\otimes\mathfrak{sp}(t,\mathbb{C})$, where $n=st,s\geq 2,t\geq 3$, or $s=2,t=4$.
@@
    $\mathfrak{h}^{c}$ is irreducible and simple.
We begin with Dynkin’s classification of simple complex $\mathfrak{h}^{c}$ such that the embedding $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ is an irreducible representation of $\mathfrak{h}^{c}$ in one of the classical Lie algebras $\mathfrak{g}^{c}=\mathfrak{sl}(V)$,$\mathfrak{so}(V),\mathfrak{sp}(V)$ for some complex vector space $V$. It says that essentially, maximal and irreducible classes coincide with several exceptions.
Theorem 10 (Dynkin, [30]).
Let $R:\mathfrak{h}\rightarrow\mathfrak{gl}(V)$ be a nontrivial irreducible representation of a simple complex Lie algebra $\mathfrak{h}$. If there are no $R(\mathfrak{h})$-invariant bilinear form on $V$, then $R(\mathfrak{h})$ is maximal in $\mathfrak{sl}(V)$. If $R(\mathfrak{h})$ preserves some orthogonal or symplectic form, then $R(\mathfrak{h})$ is maximal in $\mathfrak{so}(V)$ or $\mathfrak{sp}(V)$. The only exceptions are given in Table 1 of [14] (or Table 7 of [30]).
Remark 2.
Note that exception means "non-maximal" in this theorem.
Theorem 11 (Dynkin, [14], Theorem 2.3).
The list of all triples of irreducible inclusions $\mathfrak{h}^{c}\subset\tilde{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$ of simple Lie algebras $\mathfrak{h}^{c},\tilde{\mathfrak{h}}^{c},\mathfrak{g}^{c}$ is given by Table 5 in [14].
Comparing Theorems 9 and 10 we get the following corollary.
Corollary 5.
Any irreducible simple classical subalgebra $\mathfrak{h}^{c}$ of classical Lie algebra $\mathfrak{g}^{c}$ is either maximal, or is contained in Table 1 of [14] (or, equivalently, in Table 7 of [30]), or in Table 5 in [14]. If the representation of $\mathfrak{h}^{c}$ in $\mathfrak{g}^{c}$ is fundamental, then it is contained in Tables 1 and 2 of [7].
Remark 3.
Reading the formulations in [14] one may get an impression that Table 1 contains all the possibilities for irreducible simple classical subalgebras. However Table 5 [14] is necessary, since Table 1 of [14] and tables of fundamental representations are organized in the way which does not cover several possible triples. This is because, for example, $\mathfrak{sp}(n,\mathbb{C})$ is an irreducible representation of minimal dimension $2n$, which belongs to the table of symplectic irreducible representations, but also $\mathfrak{sp}(n,\mathbb{C})$ is contained in $\mathfrak{sl}(2n,\mathbb{C})$. Such cases are covered by Theorem 11.
5.1.2 Using Dynkin’s classification
We summarize the procedure for classification of standard Clifford-Klein forms based on the previous results as follows.
@@ 5.1.2 Using Dynkin’s classification
The classification of possible standard Clifford-Klein forms of classical simple Lie algebras amounts to finding these as real forms of the following classes of triples $(\mathfrak{g}^{c},\mathfrak{h}^{c},\mathfrak{l}^{c}$, where $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ are simple subalgebras (not necessarily proper) of the following lists of subalgebras:
  1. 1.
    the class of Levi subalgebras of parabolic subalgebras,
  2. 2.
    the class of all maximal irreducible simple subalgebras in classical $\mathfrak{g}^{c}$, or belong to Tables 1 and 5 of [14],
  3. 3.
    $\mathfrak{h}^{c}$ or $\mathfrak{l}^{c}$ is simple, reducible and is contained in one of the following inclusions:
    $\mathfrak{h}^{c}\subset\mathfrak{so}(k,\mathbb{C})\oplus\mathfrak{so}(n-k)\subset\mathfrak{so}(n,\mathbb{C}),$
    $\mathfrak{h}^{c}\subset\mathfrak{sp}(k,\mathbb{C})\oplus\mathfrak{sp}(n-k,\mathbb{C})\subset\mathfrak{sp}(n,\mathbb{C}),0<k<n,$
    $\mathfrak{h}^{c}\subset\mathfrak{so}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})\subset\mathfrak{sl}(st,\mathbb{C}),\ s,t\geq 2,$
    $\mathfrak{h}^{c}\subset\mathfrak{so}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})\subset\mathfrak{so}(st,\mathbb{C}),\ n=st,s,t\geq 3,s,t\neq 4,$
    $\mathfrak{h}^{c}\subset\mathfrak{sp}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})\subset\mathfrak{so}(st,\mathbb{C}),\ s,t\geq 2,$
    $\mathfrak{h}^{c}\subset\mathfrak{sp}(s,\mathbb{C})\otimes\mathfrak{sp}(t,\mathbb{C})\subset\mathfrak{sp}(st,\mathbb{C}),\ s\geq 2,t\geq 3,\ or\ s=2,t=4,$
    (and the same inclusions may hold for $\mathfrak{l}^{c}$),
  4. 4.
    a class of simple reducible subalgebras such that they are contained in some simple irreducible subalgebra $\tilde{\mathfrak{h}}^{c}$ of $\mathfrak{g}^{c}$ (with $\tilde{\mathfrak{h}}^{c}$ necessarily satisfying assumptions of 1)).
  5. 5.
    exceptional simple Lie subalgebras with the following simplest representations $\mathfrak{g}_{2}\subset\mathfrak{so}(7,\mathbb{C}),$ $\mathfrak{f}_{4}\subset\mathfrak{so}(26,\mathbb{C}),$ $\mathfrak{e}_{6}\subset\mathfrak{sl}(27,\mathbb{C}),$ $\mathfrak{e}_{7}\subset\mathfrak{sp}(56,\mathbb{C}),$ $\mathfrak{e}_{8}\subset\mathfrak{so}(248,\mathbb{C}).$
5.1.3 Non-generic case $\mathfrak{h}^{c}\subset B_{n}\subset D_{n+1}$
This case is in a sense "non-generic" and we consider it first. Note that it will be used in other considerations as well. We will essentially use the following.
Theorem 12 ([21], Theorem 1.5).
  • $\mathfrak{a}_{g^{\prime}}\subseteq\mathfrak{a}_{h}$,
  • $d(G^{\prime})>d(H)$.
Proposition 6.
Let
$\mathfrak{h}^{c}\subset B_{n}\subset D_{n+1}$
where $\mathfrak{h}^{c}$ is a regular simple subalgebra of $B_{n}$, and $B_{n}$ is a subalgebra of all fixed points of an outer automorphism of $D_{n+1}$. The pair $(\mathfrak{g},\mathfrak{h})$ cannot generate a standard compact Clifford-Klein form such that $\mathfrak{g}\neq\mathfrak{h}+\mathfrak{l}$ (for some $\mathfrak{l}\subset\mathfrak{g}$).
@@ 5.1.3 Non-generic case $\mathfrak{h}^{c}\subset B_{n}\subset D_{n+1}$
Choose a Cartan subalgebra $\mathfrak{c}\subset D_{n+1}$ such that $\mathfrak{c}\cap\mathfrak{h}^{c}$ is a Cartan subalgebra of $\mathfrak{h}^{c}$. The relation between roots and root vectors for the embedding $B_{n}\subset D_{n+1}$ is well known. Following [16], p. 507, we write $H_{1},...,H_{n+1}$ for the simple co-roots and the root vectors corresponding to them as $X_{1},...,X_{n+1},$. The opposite root vectors are denoted by $Y_{1},...,Y_{n+1}$. The following vectors determine the embedding of $B_{n}$ into $D_{n+1}$:
@@
Since any subalgebra $\mathfrak{h}^{c}$ which is regular in $B_{n}$ is obtained from the extended Dynkin diagram of $B_{n}$ by deleting some vertices, we can write all the possibilities for the simple roots and root vectors of $\mathfrak{h}^{c}$ in terms of the simple roots, or corresponding co-roots and root vectors expressed via the data for the type $D_{n+1}$. The extended Dynkin diagram for $B_{n}$ is given by vertices
@@
where $\alpha_{0}$ is the minimal root of $B_{n}$. Looking at the explicit expressions for maximal roots (for example, in [9]) one obtains the expression for each root of this diagram expressed via roots in of $D_{n+1}$:
@@
5.1.4 Completion of proof for $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ simple
Proposition 7.
If $\mathfrak{h}^{c}$ is a simple non-regular subalgebra in a simple Lie algebra $\mathfrak{g}^{c}$ of classical type, then
$\operatorname{rank}\mathfrak{h}^{c}\leq\frac{1}{2}(\operatorname{rank}\mathfrak{g}^{c}+1)$
with the following exceptions:

### v3

@@ 4.4.2 Construction of $Y$
For any $g\in D_{2}^{\sigma}$ the embedding $\operatorname{Ad}_{g}(\mathfrak{h})\hookrightarrow\mathfrak{g}$ is good. Thus, Theorem 4 can be applied to this embedding.
@@
Since $\mathfrak{d}_{2}^{\sigma}$ centralizes $iX$ and $\mathfrak{a}_{h}$ one derives that for any $g$ in the maximal compact subgroup of $D_{2}^{\sigma}$ the subalgebra $Ad_{g}\mathfrak{h}$ is well embedded (because $\textrm{Ad}\mathfrak{n}_{h}^{\pm}\subset\mathfrak{n}^{\pm},$ $\textrm{Ad}\mathfrak{a}_{h}=\mathfrak{a}_{h}$ and $\theta(\textrm{Ad}_{g}\mathfrak{h})=\textrm{Ad}_{g}\mathfrak{h}$). ∎
Lemma 9.
For any $\alpha\in\Phi$ vector
$0\neq\tilde{Y}:=x_{\alpha}+\sigma(x_{\alpha})+\theta(x_{\alpha}+\sigma(x_{\alpha}))$
does not belong to $\mathfrak{h}$, and for any $g\in D_{2}^{\sigma}$ vector $Y=\operatorname{Ad}_{g}(\tilde{Y})$ does not belong to $\operatorname{Ad}_{g}(\mathfrak{h})$.
Proof.
Since $x_{\alpha}+\sigma(x_{\alpha})\in\mathfrak{n}$, there exist $H\in\mathfrak{n}_{h}$ and $L\in\mathfrak{n}_{l}$ such that $x_{\alpha}+\sigma(x_{\alpha})=H+L$ and so $\tilde{Y}=H+\theta(H)+L+\theta(L).$ If $\tilde{Y}\in\mathfrak{h}$ then $L+\theta(L)\in\mathfrak{h}.$ But
@@
and thus $L+\theta(L)=H_{1}+\theta(H_{1})$ for some $H_{1}\in\mathfrak{n}_{h}.$ As $\mathfrak{n}\cap\theta(\mathfrak{n})=\{0\}$ therefore $L\in\mathfrak{n}_{h}$ and so $L=0.$ Thus $x_{\alpha}+\sigma(x_{\alpha})\in\mathfrak{n}_{h}$ and $\alpha|_{\mathfrak{a}_{h}}=0,$ a contradiction. The second claim is clear. ∎
Now we prove the main proposition which yields $Y$ as in the title of Subsection 4.4.
Proposition 2.
There exists a non-zero $Y\in\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma},$ $Y\notin\mathfrak{m}_{0}$ which centralizes $\mathfrak{h}.$
@@
For the convenience of the reader we first summarize facts needed in the course of the proof.
  1. 1.
    For any semisimple $\mathfrak{g}$ with a Cartan involution $\theta$ and any $\theta$-invariant Cartan subalgebra $\mathfrak{b}$ one has the decomposition
    $\mathfrak{b}=\mathfrak{b}_{+}\oplus\mathfrak{b}_{-}.$
    By an abuse of language we will call $\mathfrak{b}_{+}$ the compact part of $\mathfrak{b}$ and $\mathfrak{b}_{-}$ non-compact part.
  2. 2.
    By Lemma 5 $\mathfrak{c}^{\sigma}$ is a $\theta$-invariant subalgebra of $\mathfrak{d}^{\sigma}$ and, hence,
    $\mathfrak{c}^{\sigma}=\mathfrak{c}^{\sigma}_{+}\oplus\mathfrak{c}^{\sigma}_{-}$
    and by Corollary 2
    $\mathfrak{c}_{-}^{\sigma}=\mathfrak{a}_{h},\mathfrak{c}^{\sigma}=\mathfrak{c}_{+}^{\sigma}\oplus\mathfrak{a}_{h}.$
  3. 3.
    By Corollary 3
    $\mathfrak{c}^{\sigma}_{+}=\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma}.$
  4. 4.
    It is easy to see that any two compact parts of any two $\theta$-invariant Cartan subalgebras (with compact parts of the same dimension) in any semisimple real Lie algebra $\mathfrak{g}$ are conjugate, and, in particular, any two Cartan subalgebras in $\mathfrak{d}_{2}^{\sigma}$ are conjugate by an element $\operatorname{Ad}_{g},g\in D_{2}^{\sigma}$. Therefore, any compact part of any Cartan subalgebra in $\mathfrak{d}_{2}^{\sigma}$ is conjugate to $\mathfrak{d}_{2}^{\sigma}\cap\mathfrak{c}^{\sigma}$.
Now we complete the proof using 1)-4) above. Take
$\tilde{Y}=x_{\alpha}+\sigma(x_{\alpha})+\theta(x_{\alpha}+\sigma(x_{\alpha}),\alpha\in\Phi.$
By formula $(5)$, $\tilde{Y}\in\mathfrak{k}_{d}\subset\mathfrak{d}_{2}^{\sigma}$. Since $\mathfrak{k}_{d}$ is compact, $\tilde{Y}$ belongs to some compact part of some Cartan subalgebra in $\mathfrak{d}_{2}^{\sigma}$. It follows that $\tilde{Y}$ is conjugate to element $Y\in\mathfrak{c}^{\sigma}_{+}=\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma}$. Thus, $Y=\operatorname{Ad}_{g}(\tilde{Y})$ for some $g\in D_{2}^{\sigma}$. By Lemma 8 we can replace the pair $(\mathfrak{g},\mathfrak{h})$ by $(\mathfrak{g},\operatorname{Ad}_{g}(\mathfrak{h})$ and $\tilde{Y}$ by $Y\in\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma}$. By construction, $Y\not\in\mathfrak{m}_{0}$ and by Lemma 9 $Y\not\in\operatorname{Ad}_{g}(\mathfrak{h})$, hence one can also choose $Y\perp\operatorname{Ad}_{g}(\mathfrak{h})$. Since $\mathfrak{h}^{c}$ is regular, this is the same as $Y\in\mathfrak{z}_{\mathfrak{g}}(\operatorname{Ad}_{g}(\mathfrak{h}))$. Now, by some abuse of notation we write this as in the title.
@@ 5 Completion of proof of Theorem 2: non-regular case
In view of Theorem 7 we do not assume that the embeddings $\mathfrak{h}^{c},\mathfrak{l}^{c}\hookrightarrow\mathfrak{g}^{c}$ are regular. Instead, we assume that $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ are simple.
The proof is based on a careful consideration of possibilities for the real forms of the embeddings $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$, therefore, we postpone the calculations to the very end of the article. In this section we explain the main steps of the proof.
5.1 Terminology and notation
We begin by introducing some convenient (for us) terminology. Since $\mathfrak{g}^{c}$ is a classical Lie algebra, that is, $\mathfrak{g}^{c}$ can be $\mathfrak{sl}(V),\mathfrak{so}(V)$, or $\mathfrak{sp}(V)$, we get linear representations $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ in the complex vector space $V$ and call them linear if $\mathfrak{g}^{c}=\mathfrak{sl}(V)$, orthogonal, if $\mathfrak{g}^{c}=\mathfrak{so}(V)$ and symplectic, if $\mathfrak{g}^{c}=\mathfrak{sp}(V).$
We single out tensor representations, that is, the representations of the form
@@ 5.1 Terminology and notation
    $\mathfrak{h}^{c}=\mathfrak{so}(V_{1})\oplus\mathfrak{so}(V_{2})\hookrightarrow\mathfrak{g}^{c}=\mathfrak{sl}(V_{1}\otimes V_{2})$,
@@
    $\mathfrak{h}^{c}=\mathfrak{so}(V_{1})\oplus\mathfrak{so}(V_{2})\hookrightarrow\mathfrak{g^{c}}=\mathfrak{so}(V_{1}\otimes V_{2}$,
  3. 3.
    $\mathfrak{h}^{c}=\mathfrak{sp}(V_{1})\oplus\mathfrak{sp}(V_{2})\hookrightarrow\mathfrak{g}^{c}=\mathfrak{sp}(V_{1}\otimes V_{2}).$
The third type of subalgebras we need has the form
  1. 1.
    $\mathfrak{h}^{c}=\mathfrak{so}(k,\mathbb{C})\oplus\mathfrak{so}(n-k,\mathbb{C})\hookrightarrow\mathfrak{g}^{c}=\mathfrak{so}(n,\mathbb{C})$,
  2. 2.
    $\mathfrak{h}^{c}=\mathfrak{sp}(k,\mathbb{C})\oplus\mathfrak{sp}(n-k,\mathbb{C})\hookrightarrow\mathfrak{g}^{c}=\mathfrak{sp}(n,\mathbb{C}).$
We will call such subalgebras symmetric (they are fixed points of involutive automorphisms of $\mathfrak{g}^{c}$, because they are maximal in $\mathfrak{g}^{c}$ and Dynkin’s classification of such subalgebras is up to conjugation). Note that we do not consider all possible tensor representations or symmetric subalgebras, but only the listed ones. Alternatively we refer to Table 6, where maximal subalgebras of maximal rank are presented.
@@
5.2 Maximal subalgebras of classical complex simple Lie algebras
In this subsection we summarize Dynkin’s results on maximal subalgebras in classical Lie algebras in [13] and [33] which we formulate as Theorem 8 and Theorem 9.
Theorem 8 ([33]).
All maximal proper subalgebras in classical complex Lie algebras fall into the classes below, up to conjugation in $\mathfrak{g}^{c}$:
I: Reducible
  • maximal parabolic (I.1),
  • symmetric (I.2)
II: Irreducible
  • simple (II.1)
  • tensor representations (II.2)
Not all irreducible simple subalgebras are maximal, however, we have the following ( $\omega_{r}$ is 00 if the representation is linear, $1$ if it is orthogonal, and $(-1)$ if symplectic).
Theorem 9 ([13]).
Irreducible simple subalgebras in classical Lie algebras are maximal, with exceptions given by the following two tables.
Irreducible triples
Inclusion 1 Inclusion 2 $\dim$ of the representation
$B_{n}\subset A_{2n}$ $A_{2n}\subset\tilde{\mathfrak{g}}^{c}$ $\binom{2n+1}{k}$
$D_{n}\subset A_{2n-1}$ $A_{2n-1}\subset\tilde{\mathfrak{g}}^{c}$ $\binom{2n}{k}$
Table 2:
Non-maximal irreducible subalgebras
Type of inclusion dimension $\omega_{r}$
$A_{n}\subset A_{l}$ $3\binom{n+2}{4}$ 00
$A_{n}\subset A_{l}$ $3\binom{n+3}{4}$ 00
Table 3:
5.3 Steps of proof of Theorem 2
Proposition 3.
If $\mathfrak{h}^{c}$ is a simple non-regular subalgebra in $\mathfrak{g}^{c}$ then the inequality
$\operatorname{rank}\mathfrak{h}^{c}\leq\frac{1}{2}(\operatorname{rank}\mathfrak{g}^{c}+1)$ (6)
holds for all pairs $(\mathfrak{g}^{c},\mathfrak{h}^{c})$ except the following:
@@ 5.3 Steps of proof of Theorem 2
    $\mathfrak{h}^{c}=B_{p-1}\subset D_{p}$,
@@
    $\mathfrak{h}^{c}\subset B_{p-1}\subset D_{p}\subset B_{n}$, where $D_{p}\subset B_{n}$ is regular,
@@
    $\mathfrak{h}^{c}\subset B_{p-1}\subset D_{p}\subset D_{n}$, where $D_{p}\subset D_{n}$ is regular.
Proof.
Consider first the case when $\mathfrak{h}^{c}$ is irreducible. If $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ is a fundamental representation, Table 5 shows that (6) does hold (with exceptions 1)-3)). Weyl dimension formula shows that the same is valid for any irreducible representation. If $\mathfrak{h}^{c}$ is reducible, one uses Theorem 8 and obtains an inclusion $\mathfrak{h}^{c}\hookrightarrow\tilde{\mathfrak{h}}^{c}_{1}$, where $\tilde{\mathfrak{h}}^{c}_{1}$ denotes the simple summand of some maximal subalgebra $\tilde{\mathfrak{h}}^{c}$ (one can also consult Table 6). Since $\operatorname{rank}\tilde{\mathfrak{h}}^{c}_{1}<\operatorname{rank}\mathfrak{g}^{c}$, one can argue as follows. If $\mathfrak{h}^{c}\subset\tilde{\mathfrak{h}}^{c}$ where $\tilde{\mathfrak{h}}^{c}$ is parabolic, then $\mathfrak{h}^{c}$ is not regular in $\tilde{\mathfrak{h}}^{c}$ and one can argue by induction. If $\tilde{\mathfrak{h}}^{c}$ is symmetric, then the following two possibilities may occur: either $\mathfrak{h}^{c}$ is still regular in $\mathfrak{g}^{c}$, and we apply already proved part of Theorem 2, or it is one of the exceptions 1)-3). If $\mathfrak{h}^{c}\subset\tilde{\mathfrak{h}}^{c}$ is a tensor representation, then the difference between $\operatorname{rank}\mathfrak{h}^{c}$ and $\mathfrak{g}^{c}$ is even bigger. ∎
@@
If $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ and $\mathfrak{l}^{c}\hookrightarrow\mathfrak{g}^{c}$ are embeddings of simple and non-regular $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$, then for any real forms $\mathfrak{h}\hookrightarrow\mathfrak{g}$ and $\mathfrak{l}\hookrightarrow\mathfrak{g}$ either the inequality $d(\mathfrak{g})>d(\mathfrak{h})+d(\mathfrak{l})$ holds (and hence the triple $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ cannot yield compact Clifford-Klein forms), or $\mathfrak{h}^{c}=B_{p-1}$ is embedded into $\mathfrak{g}$ as in cases 1)-3) in Proposition 3. However, none of the embeddings 1)-3) above can yield compact Clifford-Klein forms.
@@
The proof of the first part of this Proposition is a rather complicated case-by case calculation of non-compact dimensions and we postpone it until subsection 5.3.1. Here we settle the cases of embeddings 1)-3) and explain the method of computing $d(\mathfrak{h})$ and $d(\mathfrak{l})$ for the real forms of simple irreducible subalgebras which are not of types 1)-3). Cases 1)-3). First, consider the embedding
$\mathfrak{h}^{c}\subset B_{n}\subset D_{n+1},$
where $\mathfrak{h}^{c}$ is a regular simple subalgebra of $B_{n}$, and $B_{n}$ is a subalgebra of all fixed points of an outer automorphism of $D_{n+1}$. The pair $(\mathfrak{g},\mathfrak{h})$ cannot generate a standard compact Clifford-Klein form such that $\mathfrak{g}\neq\mathfrak{h}+\mathfrak{l}$ (for some $\mathfrak{l}\subset\mathfrak{g}$). Choose a Cartan subalgebra $\mathfrak{c}\subset D_{n+1}$ such that $\mathfrak{c}\cap\mathfrak{h}^{c}$ is a Cartan subalgebra of $\mathfrak{h}^{c}$. The relation between roots and root vectors for the embedding $B_{n}\subset D_{n+1}$ is well known. Following [15], p. 507, we write $H_{1},...,H_{n+1}$ for the simple coroots and the root vectors corresponding to them as $X_{1},...,X_{n+1},$. The opposite root vectors are denoted by $Y_{1},...,Y_{n+1}$. The following vectors determine the embedding of $B_{n}$ into $D_{n+1}$:
@@
Since any subalgebra $\mathfrak{h}^{c}$ which is regular in $B_{n}$ is obtained from the extended Dynkin diagram of $B_{n}$ by deleting some vertices, we can write all the possibilities for the simple roots and root vectors of $\mathfrak{h}^{c}$ in terms of the simple roots, or corresponding coroots and root vectors expressed via the data for the type $D_{n+1}$. The extended Dynkin diagram for $B_{n}$ is given by vertices
@@
Now we are ready to explain the method of calculating the non-compact dimensions for all embeddings of simple non-regular Lie subalgebras.